OpenStudy (anonymous):

The position of an object at time t is given by s(t) = -8 - 9t. Find the instantaneous velocity at t = 1 by finding the derivative.

OpenStudy (amistre64):

we sem to be getting this question alot today ... its basically asking you what the slope of the line is.

OpenStudy (amistre64):

\[\lim\frac{m(x+h)-mx}{h}\] \[\lim\frac{mx+mh-mx}{h}\] \[\lim\frac{mh}{h}\] \[\lim~m=m\]